# Scalable Fine-tuning from Multiple Data Sources: A First-Order Approximation Approach

### v2

@@ 4.1 Experimental setup
Our algorithm is broadly applicable to estimating LM fine-tuning performances. In this section, we focus on instruction tuning and chain-of-thought fine-tuning. For the former, given a set of source tasks and a target task, we aim to select source tasks relevant to the target task. For the latter, we fine-tune an LM to generate chain-of-thought reasoning steps and the answer to a question. Several explanations are possible, while some of them are incorrect. We aim to select explanations pertinent to the reasoning task using subset selection.
For instruction tuning, we use three datasets, including FLAN V2 with 1,691 tasks, Chain of Thoughts with 18 tasks, and Alpaca with 38 tasks. These datasets encompass over 150 task categories. See Appendix D for the statistics.
@@
4.2 Approximating language model fine-tuning losses
@@
More specifically, we obtain $f(S)$ by fine-tuning a pretrained LM on the samples of $S$ along with the samples from the target task. We measure computation cost through the total number of floating-point operations (FLOPs) and the number of GPU hours (measured on a desktop with three Nvidia RTX6000 GPUs).
We evaluate the relative distance on both Alpaca and StrategyQA using five different LMs listed in Table 3. Due to computation constraints, for models with more than 1 billion parameters, we sample $100$ subsets to estimate approximation error for random ensemble.
@@
Figure 3: Illustration of the tradeoff between the number of FLOPs (computation cost) and test error rate, measured on our methods and six baseline methods. Our approach delivers comparable downstream performance to conventional subset selection methods, while using less than 0.5% of total computation. Furthermore, our approach achieves a 3.8% performance improvement over existing data selection methods using comparable computation costs. In particular, MTL refers to the standard approach of training one model on the data of all tasks, and we can see that our approach can indeed reduce the error rates compared to MTL. The comparison of GPU hours is qualitatively similar and can be found in Appendix D.3.
@@
4.3 Downstream evaluation for fine-tuning
Baselines: We compare our algorithms to forward selection (FS) and random ensemble (RE) with full fine-tuning. Additionally, we compare with four baseline methods, including fine-tuning a single model on all tasks with LoRA (MTL), Data Selection with Importance Resampling (DSIR) Xie et al. (2023), Data-Efficient Fine-Tuning with cross-task nearest neighbors clustering based on feature similarity (DEFT) Ivison et al. (2023), and Low-rank gradiEnt Similarity Search (LESS) Xia et al. (2024). For each baseline, the performance is reported based on fine-tuning a pretrained LM with the same number of trainable parameters on the selected subset of tasks.
Evaluation metrics: For instruction tuning, we follow the protocol of Wang et al. (2023a) to evaluate the toxicity and truthfulness of fine-tuned models in ToxiGen Hartvigsen et al. (2022) and TruthfulQA Lin et al. (2022), respectively. For ToxiGen, we measure the percentage of toxic outputs generated by the model. For TruthfulQA, we evaluate model accuracy for identifying truthful statements. For StrategyQA and CommonsenseQA, we measure the accuracy of the generated answers.
4.3.1 Results for instruction tuning
@@
4.3.2 Results for chain-of-thought fine-tuning
Next, we report the results on chain-of-thought fine-tuning, illustrated in Figure 3(b), tested on FLAN-T5-Base. In a nutshell, our findings remain consistent with those of Section 4.3.1.
First, GradEx-FS outperforms directly testing the pretrained model and feature transfer by 7% and 16%, respectively. GradEx-FS also improves the accuracy by 5% over MTL. Second, even with the estimation, we see similar downstream performance to forward selection and random ensemble (with true fine-tuning) while using only 3% and 0.5% of the total computation costs. Lastly, both GradEx-FS and GradEx-RE can outperform the four baseline selection methods by 2.4% and 3.5% (on average), respectively.
As for data selection, we first cluster the samples into $n$ groups. We vary $n$ between $50$, $100$, $200$, and $400$. We also evaluate our approach without the clustering step. We find that using $100$ groups yields the best performance and indeed outperforms not using clustering by 1.4%. Therefore, we set the number of groups as $100$ in the chain-of-though fine-tuning.
@@ 5 Related Work
Parameter-efficient fine-tuning: One influential line of work has sought to design adapters, which are small modules injected into the intermediate layers of a deep neural network. With adapters, only a small fraction of the entire network has to change inside the adapters, and this approach has found applications in many settings such as text classification Howard and Ruder (2018), text transfer Pfeiffer et al. (2021), and cross-lingual transfer such as named entity recognition and commonsense reasoning Bapna and Firat (2019). Another different approach is to use LoRA Hu et al. (2021), which constrains the fine-tuning region inside a low-rank subspace, greatly improving the parameter efficiency of fine-tuning. Both LoRA and follow-up works such as QLoRA Dettmers et al. (2023) and ReLoRA Lialin et al. (2024) focus on fine-tuning a single task. Mahabadi et al. (2021) studies parameter-efficient multitask fine-tuning via shared hypernetworks. Our work can be viewed as expanding these methods to multitask learning (MTL). We believe this connection between MTL and fine-tuning would be worth further exploration in future work.
@@
Data modeling: There is a growing line of work on understanding the role of individual samples in deep networks and large models. Kozareva et al. (2011) propose a centrality and regularization-based algorithm by leveraging graph structures to explore the relationship between multiple instances in a semantic network. Wettig et al. (2024) select pre-training data by training a rater model to evaluate four data quality criteria that align with human intuitions. Thrush et al. (2024) select data based on the losses and perplexities of existing LLMs on pretraining texts. In contrast, this work focuses on task-relatedness in fine-tuning language models. This is a simpler problem than pre-training, as the number of tasks is smaller. A potential adaptation could be splitting pre-training into multiple phases and applying our techniques to develop a curriculum for pre-training.
Our approach is also related to the work of Ilyas et al. (2022) and Park et al. (2023) on using random sampling for data modeling and attribution. The difference is that we utilize the linearization property of LLMs for fine-tuning close to the initialization. A closely related work by Li et al. (2023b) introduces a surrogate modeling approach to identify subsets of source tasks that benefit a target task. They show that a linear surrogate model is particularly helpful for identifying negative transfers, including higher-order negative transfers from a subset of source tasks to the target task. This higher-order transfer corresponds to a boosting procedure that turns out to be very helpful in practice Li et al. (2023a). In addition, this surrogate modeling approach can also be used in data augmentation for finding tree-structured compositions Li et al. (2024a). It would be very interesting to understand the underlying mathematical structure of LLMs that enables such geometric properties in practice Zhang (2023).
@@ 6 Conclusion
This paper introduced a novel method for estimating LM fine-tuning performances with a first-order approximation approach. The method can significantly accelerate conventional subset selection, thus unlocking their applications to task/data selection for fine-tuning. Evaluation across numerous datasets and LMs demonstrates the benefit of GradEx when applied to subset selection.
Acknowlegement: Thanks to the anonymous referees and the action editor for their constructive feedback. The work of D. Li, Z. Zhang, and H. Zhang is partly supported by NSF award IIS-2412008. D. Li was also partially funded by a PhD fellowship from JPMorgan Chase & Co. Any views or opinions expressed herein are solely those of the authors listed, and may differ from the views and opinions expressed by JPMorgan Chase & Co. or its Affiliates.
Appendix A Mathematical Notations
We provide a list of symbols and their meaning used in the paper for reference below:
  • $S$: A subset of $\{1,2,\dots,n\}$.
@@ Appendix B Omitted Materials from Section 2
$f(S)$ is not monotone: We find that $f(S)$ is not monotone, in the sense that if we add one helper task into $S$, this does not necessarily improve the outcome. Concretely, we start with an initial set $S$. It contains a target task $t$ corresponding to the instructions of “arrange” in Alpaca. Then, we keep adding a “helping” task ($i$) to $S$, if $f(\{i,t\})<f(\{t\})$ (reducing the loss). According to our experiment, once more than two tasks are added, $f(S)$ increases gradually, indicating that adding more tasks can worsen the performance of the target task, and these are all “helpful” tasks (in the sense of pairwise transfer).
$f(S)$ is not submodular:
Appendix C Ommited Materials for Section 3
Meta-training improves approximation quality: In Section 3, we observe that first-order expansion near a meta-initialization fine-tuned on all tasks provides an accurate approximation. Next, we evaluate whether this approximation can be achieved without the meta-training step. We measure the first-order Taylor expansion on the pretrained GPT-Neo-1.3B and TinyLlama-1.1B model. We observe that the approximation from the pretrained initialization incurs within 7% error. In contrast, using meta-initialization is significantly better. The approximation achieves less than 1% error. We report the results in Table 4.
@@ Appendix C Ommited Materials for Section 3
While using the second-order approximation can further reduce estimation error, it requires computing Hessian-gradient products, which is computationally expensive. Thus, we focus on the first-order approximation in this paper.
Dimension reduction:
@@
Application to multi-classification tasks:
@@
Application to generative tasks:
@@
Table 5: Detailed statistics about all the datasets used in the experiments.
Dataset # Tasks Avg size Category Source
@@
Table 6: List of hyper-parameters used in the experiments.
@@
Runtime complexity.
@@
Forward selection: This algorithm performs greedy subset selection. It starts with an empty set and iterates up to $n$ times to select the optimal task to combine with the current subset. At the $i$-th iteration, it fine-tunes one model on the current subset combined with each of the remaining $n-i+1$ unselected tasks. Consequently, the total number of forward passes is at most $\sum_{i=1}^{n}(n-i+1)\cdot i=\frac{1}{6}n(n+1)(n+2)$.
@@
DSIR Xie et al. (2023): This algorithm assesses the $n$-gram features of every data sample without requiring model forward passes. It trains a model on a selected subset of tasks. The total number of forward passes is at most $n$.
DEFT Xia et al. (2024): These two algorithms follow a similar procedure. First, train a model on all given tasks. Then, compute the similarity score between feature representations (or gradients) of training and test samples. Lastly, select tasks based on the scores and train a model on the selected subset of tasks. This process requires a total of $3n$ forward passes.
GradEx: Our algorithm requires a comparable number of forward passes as DEFT. First, we train a model on all task data and project the gradients of each data sample. Then, instead of computing gradient similarities, we estimate model fine-tuning performances by solving logistic regression on the projected gradients. Finally, we select a subset of tasks based on these estimated performances and train a model on the selected subset. This also requires a total of $3n$ forward passes.
Appendix D Additional Experiments
D.1 Setup
@@
In our experiments, we fine-tune the models using the AdamW optimizer. We fine-tune the entire model for smaller models, including BERT-Base, RoBERTa-Base, Pythia-70M, and GPT-2. For other models, we use LoRA for fine-tuning. The training hyper-parameters used for each experiment are described in Table 6.
For ToxiGen, we measure the percentage of toxic generations of the model. First, the model is prompted to produce toxic languages using human-designed hateful prompts; then, the generations are classified by a toxic content detection model. For TruthfulQA, we evaluate the model’s accuracy in detecting truthful statements. Given a question and five answer choices, we measure the accuracy of the model in assigning the highest probability to the correct answer.
D.2 Omitted results
In Table 7, we report the complete results corresponding to Figure 3.
D.3 Comparison of GPU hours
We also compare the GPU hours of each method on StrategyQA. We observe that GradEx-FS uses 10.6 GPU hours and GradEx-RE uses 14.7 GPU hours. Corroborating with the FLOPs comparison results, our algorithm reduces the GPU hours of classic selection methods by 10$\times$. Moreover, our algorithm uses comparable GPU hours to existing selection baselines, with LESS and DEFT taking 10.2 GPU hours. The comparison has a similar trend in instruction fine-tuning datasets.
Table 7: Performance scores and computation cost (FLOPs) on four datasets. ToxiGen and TruthfulQA are used to evaluate instruction tuning. CommonsenseQA and StartegyQA are used to evaluate chain-of-thought fine-tuning. We report the averaged results and standard deviations over three random seeds for each method.

### v3

@@ 4.1 Experimental setup
Our algorithm is broadly applicable to estimating the performance of LLM fine-tuning. In this section, we focus on instruction tuning and chain-of-thought fine-tuning. For the former, given a set of source tasks and a target task, we aim to select source tasks relevant to the target task. For the latter, we fine-tune an LM to generate chain-of-thought reasoning steps and the answer to a question. Several explanations are possible, while some of them are incorrect. We aim to select explanations pertinent to the reasoning task using subset selection.
For instruction tuning, we use three datasets, including FLAN V2 with 1,691 tasks, Chain of Thoughts with 18 tasks, and Alpaca with 38 tasks. These datasets encompass over 150 task categories. See Appendix C for the statistics.
@@
4.2 Results on approximation errors
@@
More specifically, we obtain $f(S)$ by fine-tuning a pretrained LM on the samples of $S$ along with the samples from the target task. We measure computation cost in terms of the total number of floating-point operations (FLOPs) and the number of GPU hours (measured on a desktop with three Nvidia RTX 6000 GPUs).
We evaluate the relative distance on both Alpaca and StrategyQA using five different LMs listed in Table 3. Due to computation constraints, for models with more than 1 billion parameters, we sample $100$ subsets to estimate the approximation error for the random ensemble.
@@
Figure 3: Illustration of the tradeoff between the number of FLOPs (computation cost) and test error rate, measured on our methods and six baseline methods. Our approach delivers comparable downstream performance to conventional subset selection methods, while using less than 0.5% of total computation. Furthermore, our approach achieves a 3.8% performance improvement over existing data selection methods using comparable computation costs. In particular, MTL refers to the standard approach of training a single model on the data from all tasks, and we can see that our approach indeed reduces the error rates compared to MTL. The comparison of GPU hours is qualitatively similar and can be found in Appendix C.1.
@@
4.3 Results on model fine-tuning
We compare our algorithms to forward selection (FS) and random ensemble (RE) with full fine-tuning. Additionally, we compare with four baseline methods, including fine-tuning a single model on all tasks with LoRA (MTL), Data Selection with Importance Resampling (DSIR) Xie et al. (2023), Data-Efficient Fine-Tuning with cross-task nearest neighbors clustering based on feature similarity (DEFT) Ivison et al. (2023), and Low-rank gradiEnt Similarity Search (LESS) Xia et al. (2024). For each baseline, performance is reported based on fine-tuning a pre-trained model with the same number of trainable parameters on the selected subset of tasks.
For instruction tuning, we follow the protocol of Wang et al. (2023a) to evaluate the toxicity and truthfulness of fine-tuned models in ToxiGen Hartvigsen et al. (2022) and TruthfulQA Lin et al. (2022), respectively. For ToxiGen, we measure the percentage of toxic outputs generated by the model. For TruthfulQA, we evaluate model accuracy for identifying truthful statements. For StrategyQA and CommonsenseQA, we measure the accuracy of the generated answers.
Results for instruction tuning.
@@ 4.3 Results on model fine-tuning
Results for chain-of-thought fine-tuning.
Next, we report the results on chain-of-thought fine-tuning, illustrated in Figure 3(b), tested on FLAN-T5-Base. In summary, our findings are consistent with those of instruction tuning.
First, GradEx-FS outperforms direct testing of the pre-trained model and feature transfer by 7% and 16%, respectively. GradEx-FS also improves the accuracy by 5% over MTL. Second, even with estimation, we observe similar downstream performance to forward selection and random ensemble (with true fine-tuning) while using only 3% and 0.5% of the total computational costs. Lastly, both GradEx-FS and GradEx-RE can outperform the four baseline selection methods by 2.4% and 3.5% (on average), respectively.
As for data selection, we first cluster the samples into $n$ groups. We vary $n$ between $50$, $100$, $200$, and $400$. We also evaluate our approach without the clustering step. We find that using $100$ groups yields the best performance and indeed outperforms not using clustering by 1.4%. Therefore, we set the number of groups as $100$ in the chain-of-thought fine-tuning.
@@ 5 Related Work
Parameter-efficient fine-tuning: One influential line of work has sought to design adapters, which are small modules injected into the intermediate layers of a deep neural network. With adapters, only a small fraction of the entire network has to change inside the adapters, and this approach has found applications in many settings such as text classification Howard and Ruder (2018), text transfer Pfeiffer et al. (2021), and cross-lingual transfer such as named entity recognition and commonsense reasoning Bapna and Firat (2019). Another different approach is to use LoRA Hu et al. (2021), which constrains the fine-tuning region inside a low-rank subspace, greatly improving the parameter efficiency of fine-tuning. Both LoRA and follow-up works, such as QLoRA Dettmers et al. (2023) and ReLoRA Lialin et al. (2024), focus on fine-tuning a single task. Mahabadi et al. (2021) studies parameter-efficient multitask fine-tuning via shared hypernetworks. Our work can be viewed as expanding these methods to multitask learning (MTL). We believe this connection between MTL and fine-tuning would be worth further exploration in future work.
@@
Data modeling: There is a growing line of work on understanding the role of individual samples in deep networks and large models. Kozareva et al. (2011) propose a centrality and regularization-based algorithm by leveraging graph structures to explore the relationship between multiple instances in a semantic network. Wettig et al. (2024) select pre-training data by training a rater model to evaluate four data quality criteria that align with human intuitions. In contrast, this work focuses on task-relatedness in fine-tuning language models. This is a simpler problem than pre-training, as the number of tasks is smaller. A potential adaptation could be splitting pre-training into multiple phases and applying our techniques to develop a curriculum for pre-training.
Our approach is also related to the work of Ilyas et al. (2022) and Park et al. (2023) on using random sampling for data modeling and attribution. The difference is that we utilize the linearization property of LLMs for fine-tuning close to the initialization. A closely related work by Li et al. (2023b) introduces a surrogate modeling approach to identify subsets of source tasks that benefit a target task. They show that a linear surrogate model is particularly helpful for identifying negative transfers, including higher-order negative transfers from a subset of source tasks to the target task. This higher-order transfer corresponds to a boosting procedure that turns out to be very helpful in practice Li et al. (2023a). In addition, this surrogate modeling approach can also be used in data augmentation for finding tree-structured compositions Li et al. (2024a). We remark that in a follow-up paper, we provide an improved version of our algorithm for low-rank adapter tuning Li et al. (2025). Moreover, we find that for many LLMs, one does not need the meta-training stage to apply such low-rank adaption methods before using the linearization technique. It would be very interesting to understand the underlying mathematical structure of LLMs that enables such geometric properties in practice Zhang (2023).
@@ 6 Conclusion
This paper introduced a novel method for estimating LM fine-tuning performances with a first-order approximation approach. The method can significantly accelerate conventional subset selection, thus unlocking its applications to task/data selection for fine-tuning. Evaluation across numerous datasets and LMs demonstrates the benefit of GradEx when applied to subset selection.
Acknowledgments
Thanks to the anonymous referees and the action editor for their constructive feedback. The work of D. Li, Z. Zhang, and H. Zhang is partly supported by NSF award IIS-2412008. D. Li was also partially funded by a PhD fellowship from JPMorgan Chase & Co. Any views or opinions expressed herein are solely those of the authors listed, and may differ from the views and opinions expressed by JPMorgan Chase & Co. or its Affiliates.
Limitations
One limitation of our algorithm is that it requires estimating the fine-tuned model’s performance based on all the model weights. Designing techniques to estimate fine-tuning performance in a limited set of model weights or for closed-source models such as GPT-4 is an interesting open question. Our study is also restricted to fine-tuning, while there may well be other scenarios where one would like to adopt a language model, such as in-context learning or alignment with RLHF. Further exploring functional approximation techniques for these settings would be another promising avenue for future work.
Broader Implications
This paper examines the problem of fine-tuning language models given multiple data sources. While the use of language models may have potential societal consequences in the future, there are no specific ethical concerns arising from our work. Due to the technical nature of this paper, there are no direct implications of negative societal impacts.
Appendix A Omitted Materials from Section 2
Notations. We provide a list of notations and their meanings used in the paper for reference below:
  • $S$: A subset of the $n$ tasks from $\{1,2,\dots,n\}$.
@@ Appendix A Omitted Materials from Section 2
$f(S)$ is not monotone.
We find that $f(S)$ is not monotone, in the sense that if we add one helper task into $S$, this does not necessarily improve the outcome. Concretely, we start with an initial set $S$. It contains a target task $t$ corresponding to the instructions of “arrange” in Alpaca. Then, we keep adding a “helping” task ($i$) to $S$, if $f(\{i,t\})<f(\{t\})$ (reducing the loss). According to our experiment, once more than two tasks are added, $f(S)$ increases gradually, indicating that adding more tasks can worsen the performance of the target task, and these are all “helpful” tasks (in the sense of pairwise transfer).
$f(S)$ is not submodular.
@@ Appendix B Omitted Materials from Section 3
Meta-training improves approximation quality. In Section 3, we observe that the first-order expansion near a meta-initialization fine-tuned on all tasks provides an accurate approximation. Next, we evaluate whether this approximation can be achieved without the meta-training step. We measure the first-order Taylor expansion on the pretrained GPT-Neo-1.3B and TinyLlama-1.1B models. We observe that the approximation from the pretrained initialization incurs within 7% error. In contrast, using meta-initialization is significantly better. The approximation achieves less than 1% error. We report the results in Table 4.
@@
While using the second-order approximation can further reduce estimation error, it requires computing Hessian-gradient products, which are computationally expensive. Thus, we focus on the first-order approximation in this paper.
Dimension reduction.
@@
Extension to multi-classification.
@@
Extension to generative tasks.
@@
Table 5: Detailed statistics about eight datasets used in the experiments, including the number of tasks, and the link to their data source.
Dataset # Tasks Average size Category Source
@@
Table 6: List of hyperparameters for each dataset tested in the experiments, corresponding to the results that we reported in Section 4.
@@
Analysis of running time complexity.
@@
Forward selection: This algorithm employs a greedy subset selection approach. It starts with an empty set and iterates up to $n$ times to select the optimal task to combine with the current subset. At the $i$-th iteration, it fine-tunes one model on the current subset combined with each of the remaining $n-i+1$ unselected tasks. Consequently, the total number of forward passes is at most $\sum_{i=1}^{n}(n-i+1)\cdot i=\frac{1}{6}n(n+1)(n+2)$.
@@
Data selection via important resampling Xie et al. (2023): This algorithm assesses the $n$-gram features of every data sample without requiring model forward passes. It trains a model on a selected subset of tasks. The total number of forward passes is at most $n$.
Low-rank gradient similarity search Xia et al. (2024): This algorithm first trains a model on all given tasks. Then, compute the similarity score between feature representations (or gradients) of training and test samples. Lastly, select tasks based on the scores and train a model on the selected subset of tasks. This process requires a total of $3n$ forward passes.
Gradient estimation (GradEx): This algorithm requires a comparable number of forward passes as DEFT. First, we train a model on all task data and project the gradients of each data sample. Then, instead of computing gradient similarities, we estimate model fine-tuning performances by solving logistic regression on the projected gradients. Finally, we select a subset of tasks based on these estimated performances and train a model on the selected subset. This also requires a total of $3n$ forward passes.
Appendix C Omitted Experiments from Section 4
@@
In our experiments, we fine-tune the models using the AdamW optimizer. We fine-tune the entire model for smaller models, including BERT-Base, RoBERTa-Base, Pythia-70M, and GPT-2. For other models, we use LoRA for fine-tuning. The training hyperparameters used for each experiment are described in Table 6.
For ToxiGen, we measure the percentage of toxic generations of the model. First, the model is prompted to produce toxic language using human-designed hateful prompts; then, the generations are classified by a toxic content detection model. For TruthfulQA, we evaluate the model’s accuracy in detecting truthful statements. Given a question and five answer choices, we measure the accuracy of the model in assigning the highest probability to the correct answer.
C.1 Runtime-accuracy tradeoff
In Table 7, we report the complete results corresponding to the illustrations shown in the trade-off in Figure 3 of Section 4.
We also compare the GPU hours of each method on StrategyQA. We observe that GradEx-FS uses $10.6$ GPU hours and GradEx-RE uses $14.7$ GPU hours. Corroborating with the FLOPs comparison results, our algorithm reduces the GPU hours required by classic selection methods by 10×10$\times$10 ×. Moreover, our algorithm uses comparable GPU hours to existing selection baselines, with LESS and DEFT taking $10.2$ GPU hours.
We also note that the comparison has a similar trend in the instruction tuning setting.
Table 7: Performance scores and computation cost (FLOPs) on four datasets, including ToxiGen, TruthfulQA, CommonsenseQA, and StrategyQA. The ToxiGen and TruthfulQA datasets are used to evaluate the instruction tuning setting. The CommonsenseQA and StrategyQA datasets are used to evaluate the chain-of-thoughts fine-tuning setting. We report the averaged results and standard deviations over three random seeds for each method below.